OpenStudy (anonymous):

Solve 5x + y = 2 4x + y = 4

OpenStudy (anonymous):

from first equation 5x+y=2 y=2-5x in second equation 4x+y=4 4x+(2-5x)=4 4x+2-5x=4 -x+2=4 -x=4-2 -x=2 x=-2 y=2-5x =2-(5*-2)=2-(-10)=2+10 =12

OpenStudy (anonymous):

or you solve two equations together as u subtract second equation from first (5x-4x) + (y-y) =2-4 x+0=-2 x=-2 in first equation 5x+y=2 5*(-2)+y=2 -10+y=2 y=10+2 y=12